# Pritam Goswami    Adri Bhattacharya    Raja Das    Partha Sarathi Mandal

### v1

@@ 5.2.1 Correctness and Complexity
Case-II: Let $a_{j}$ gets consumed by the black hole at round $t$ while in state Backtrack. We have to show that the agent that changes the state to Gather must store the dir type message with direction component counter-clockwise (as any agent can only move in counter clockwise direction, in state Backtrack). It will be enough to show that the agent that changes its state to Gather must have changed it from state Forward (as only in this case the state changing agent stores the dir type message having a counter-clockwise direction component in its local memory). Let $t^{\prime}$ be the round when $a_{j}$ was in state Initial for the last time. We claim that $a_{i}$ and $a_{k}$ can not change to state Gather before round $t^{\prime}+3n+1$. Note that at round $t^{\prime}$ both $a_{i}$ and $a_{k}$ are at their corresponding $home$ in state Initial. Now if any one of them changes to state Gather, the earliest it can happen is at round $t^{\prime}+n+1$ when both of them are in state Forward. In this case the agent that changes to state Gather must have found a visited type message at the current node (i.e., for $a_{i}$ it is $home$ of $a_{k}$ and for $a_{k}$ it is $home$ of $a_{j}$) at round $t^{\prime}+n+1$. Note that during round $t^{\prime}$ any previous messages are erased from both $home$ of $a_{k}$ and $home$ of $a_{j}$ by $a_{k}$ and $a_{j}$, respectively and no other agent can visit and alter data at these nodes before $t^{\prime}+n+1$ round. Hence none of $a_{i}$ and $a_{k}$ finds any visited type message at their current nodes at round $t^{\prime}+n+1$. Hence, none of these agents changes their state to Gather at round $t^{\prime}+n+1$. Next they can only change their state to Gather at the round $t^{\prime}+3n+1$ when both of them (i.e., $a_{i}$ and $a_{k}$) are at their respective $home$. An agent among them changes its state to Gather at round $t^{\prime}+3n+1$ if it finds no visited type message at their current node (i.e., corresponding $home$). Note that $t^{\prime}+3n+1>t>t^{\prime}+n+1$ (as $a_{j}$ was in state Backtrack at round $t$). So, at round $t^{\prime}+n+1$ all of $a_{i},a_{j}$ and $a_{k}$ are at nodes $home$ of $a_{k},home$ of $a_{i}$ and $home$ of $a_{j}$ in state Forward and writes a visited type message in the nodes, respectively. These messages can not be altered by any agent until round $t^{\prime}+3n+1$ (by similar argument as in Case-I). Note that at round $t^{\prime}+3n+1$ both $a_{i}$ and $a_{k}$ are at their corresponding $home$ and both of them finds a visited type message at these nodes left by $a_{j}$ and $a_{i}$ respectively (at round $t^{\prime}+n+1$). So, none of them changes to state Gather even at round $t^{\prime}+3n+1$. Next they move into state Initial-Wait and wait at their $home$ until $t^{\prime}+4n$ round. Now, since no agent in state Initial-Wait can change its state to Gather2 until it meets with an agent in state Gather and all alive agent at round $t^{\prime}+3n+2$ are at state Initial-Wait all of them (i.e., $a_{i}$ and $a_{k}$) waits and changes state to Initial again at round $t^{\prime}+4n+1$. Next at round $t^{\prime}+4n+2$ both $a_{i}$ and $a_{k}$ erase all previous data at their corresponding $home$. and changes status to Forward again. Note that the next time $a_{i}$ and $a_{k}$ can change the state to gather must be at the round $t^{\prime}+5n+2$ when both of them are in state Forward, and are currently at the $home$ of $a_{k}$ and $home$ of $a_{j}$, respectively. Further, note that since $home$ of $a_{k}$ does not have any visited type message at round $t^{\prime}+5n+2$ (as $a_{k}$ erased any data at round $t^{\prime}+4n+2$ and no other agent can alter data there after $t^{\prime}+4n+2$ and before $t^{\prime}+5n+2$), so $a_{i}$ can not change to state Gather at round $t^{\prime}+5n+2$. Also observe that the visited type message at the $home$ of $a_{j}$ written by $a_{k}$ during round $t^{\prime}+3n+1$ is still there at round $t^{\prime}+5n+2$ (after $t^{\prime}+3n+1$ before $t^{\prime}+4n+2$ since $a_{j}$ is already consumed before $t^{\prime}+3n+1$, no agent is on $home$ of $a_{j}$ to erase the data, also after $t^{\prime}+4n+1$ till $t^{\prime}+5n+1$ only $a_{k}$ can visit $home$ of $a_{j}$ in state Forward and Back-Wait, but in these states it does not alter any data at the $home$ of $a_{j}$) $a_{k}$ finds it during the round $t^{\prime}+5n+2$ in state Forward and changes its state to Gather.
@@
Proof 5.16.
Suppose an agent $a_{i}$ in state Back-Wait gets a dir type message with clockwise direction component at some round $t$. This implies there exists another agent $a_{k}$ that has changed its state to Gather after storing a dir type message having direction component clockwise at some round $t^{\prime}<t$. This can only happen if $a_{k}$ was in state Backtrack at the beginning of round $t^{\prime}$. So at the beginning of round $t^{\prime}$, $a_{i}$ was also in state Backtrack and thus at round $t^{\prime}$, $a_{i}$ changes its state to Initial-Wait. Note that $a_{k}$ meets and shares dir type message with $a_{i}$ while $a_{i}$ is still at state Initial-Wait. This contradicts our assumption that $a_{i}$ gets dir message at state Back-Wait. Thus, if $a_{i}$ finds a dir type message in state Back-Wait then it must have the direction component counter-clockwise.
With similar argument we can also proof the following lemma.
@@
Let $a_{i}$, $a_{j}$ and $a_{k}$ be three agents executing the algorithm PerpExplore-Scat-Whitbrd, and suppose $a_{j}$ be the first agent to enter the black hole, while exploring the segment $Seg(a_{j})$. Let $v_{1}$ be the $home$ of $a_{j}$ and $v_{2}$ be the furthest node from $v_{1}$ along clockwise direction which is inside $Seg(a_{j})$. We define the Cautious start node to be $v_{1}$ if $a_{j}$ is consumed by the black hole during state Forward. Otherwise, if $a_{j}$ is consumed by the black hole in state Backtrack then, $v_{2}$ is defined to be the Cautious start node.
@@
Let the first agent is consumed by the black hole at some round $t>0$, then there exists a round $t^{\prime}>t$, at which the remaining alive agents reach the cautious start node and changes their state to Cautious-Leader and Cautious-Follower.
Proof 5.20.
@@
Let $a_{i}$, $a_{j}$ and $a_{k}$ be three agents, exploring $Seg(a_{i})$, $Seg(a_{j})$ and $Seg(a_{k})$, respectively, where $Seg(a_{i})\cap Seg(a_{j})=home$ of $a_{i}$, $Seg(a_{i})\cap Seg(a_{k})=home$ of $a_{k}$ and $Seg(a_{k})\cap Seg(a_{j})=home$ of $a_{j}$. Let $a_{j}$ be the agent which gets consumed by the black hole, while exploring $Seg(a_{j})$, then we have the following cases:
@@
Case-I: $a_{j}$ is consumed by the black hole while it is moving in a clockwise direction, i.e., in state Forward at round $t>0$. Note that in this case the cautious start node is the $home$ of $a_{j}$. This implies there exists a round $0<t_{0}<t<t_{0}+n$, where $a_{j}$ was in state Initial. Note that, $a_{j}$ can not write any visited type message at $home$ of $a_{i}$, as it gets consumed before reaching that node. So, at round $t_{0}+3n+1$, when $a_{i}$ is at its own $home$ in state Backtrack and checks any visited type message, it finds none exists. This triggers $a_{i}$ to change its state to Gather from Backtrack with the corresponding dir = (clockwise, NULL) (which is indeed the correct direction, refer Lemma 5.13). Further, the agent finds $a_{k}$, which is currently waiting at its own $home$ in state Initial-Wait (as it does not find any anomaly, so from Backtrack it changed to state Initial-Wait at round $t_{0}+3n+1$). So, the moment $a_{i}$ reaches the $home$ of $a_{k}$, it takes one additional round to store the dir type message at the current node and then changes its state to Gather1. On the other hand, whenever $a_{k}$ finds a dir type message is written at its current node, it also changes its state to Gather1, i.e., at round $t_{0}+3n+1<t^{\prime\prime}<t_{0}+4n$ both agents change their state to Gather1. After which, they together start moving in a clockwise direction (5.2), until they reach a node which has a home type message with the ID component, different from the IDs of $a_{i}$ and $a_{k}$. Note that this must be the $home$ of $a_{j}$, as the home type message $a_{j}$ has written at round $t_{0}$, cannot be erased by any other agent except $a_{j}$. Also $a_{j}$ can only erase this in state Initial at round $t_{0}+4n+2$ if it would have returned back, but since it is already consumed between round $t_{0}+1$ and $t_{0}+n-1$, hence this possibility never arises, so the home type message remains, when $a_{i}$ and $a_{k}$ together reaches this node. After which they change their state to Cautious-Leader and Cautious-Follower depending on their IDs.
@@
Case-II: $a_{j}$ is consumed by the black hole while it is moving in a counter-clockwise direction, i.e., in state Backtrack at round $t>0$. Note that in this case the cautious start node is the $home$ of $a_{i}$. Let $t_{0}$ be the round when $a_{j}$ was in state Initial the last time. This implies $t_{0}+2n<t<t_{0}+3n$, now by similar argument as explained in Case-II of Lemma 5.13, $a_{k}$ changes its state to Gather while storing the dir = (counter-clockwise,NULL) message, at the $home$ of $a_{j}$ from state Forward, and starts moving in a counter-clockwise direction. Note that at round $t_{0}+5n+2$ as $a_{i}$ did not find any anomaly, so it changes its state to Back-Wait from Forward. Hence, at round $t_{1}$, where $t_{1}<t_{0}+6n$, $a_{k}$ finds $a_{i}$, while $a_{i}$ is still in state Back-Wait. This triggers $a_{k}$ to change its state to Gather2 at round $t_{1}$, while updating the dir type message to (Counter-clockwise, ID’), where ID’ is the ID of $a_{i}$. Then at the same round $a_{k}$ writes the updated message at the current node (i.e., $home$ of $a_{k}$). Whenever $a_{i}$ sees this dir type message (at round $t_{1}+1$) it also changes its state to Gather2. Next in state Gather2, both start to move counter-clockwise (from round $t_{1}+2$) and continues to move until they find a home type message with ID matching the ID of the dir type message. Note that, this node is nothing but the $home$ of $a_{i}$ (as the ID component of dir type message stores the ID of $a_{i}$). Note that at round $t_{0}+4n+2$, $a_{i}$ written a home type message at its own $home$. This message can be erased only again at round $t_{0}+8n+2$ (i.e., when $a_{i}$ reaches its $home$ again in state Initial). But since $a_{i}$ changed its state to Gather2 before $t_{0}+6n+1$ it cannot move back to state Initial again. So,when $a_{k}$ and $a_{i}$ reaches $home$ of $a_{i}$, they finds the home type message. Hence, both $a_{k}$ and $a_{i}$ reach the cautious start node within $t_{0}+7n$ and further changes their states to Cautious-Leader and Cautious-Follower, depending on their IDs.
@@
Let $a_{i}$ and $a_{k}$ be the two agents which start the states Cautious-Leader and Cautious-Follower from the cautious start node, then within finite rounds of executing algorithm PerpExplore-Scat-WhitBrd, at least one agent detects the location of the black hole.
Proof 5.22.
@@
Let $a_{j}$ be the agent that has been consumed by the black hole at round $t>0$. Now there are two cases based on the state of $a_{j}$ at round $t$.
@@
Case-I: $a_{j}$ was in state Forward at round $t$. In that case the node $v_{1}=home$ of $a_{j}$ is the cautious start node. Let $v_{2}$ be the farthest node along clockwise direction in $Seg(a_{j})$ (i.e., the clockwise nearest $home$ of $a_{j}$). Now if there exists any round $t^{\prime}<t$ when $a_{j}$ was in state Backtrack then it must have started its state Backtrack from $v_{2}$ and moved counter-clockwise until $v_{1}$ while erasing all right markings from each node in $Seg(a_{j})$. So when $a_{j}$ started the state Forward at its $home$ before being consumed, all nodes in $Seg(a_{j})$ are without any right marking. This case will happen also when there are no round $t^{\prime}<t$ where $a_{j}$ was in state Backtrack, i.e., there does not exist any round before $t^{\prime}$, when $a_{j}$ was in state Forward. So before $a_{j}$ is consumed at the black hole , say $v_{b}$, at round $t$, it has marked right at all nodes starting from the next node of $v_{1}$ in clockwise direction, upto the node just before $v_{b}$ ($v_{b}$ can not be marked as before marking it the agent $a_{j}$ is consumed there). Let without loss of generality, $a_{i}$ is in state Cautious-Leader and $a_{j}$ is in state Cautious-Follower at $v_{1}$ at some round $t_{0}>t$. Then $a_{i}$ always moves ahead alone in clockwise direction to a new node $v$.Then it moves back only if sees the right marking and brings $a_{k}$ to $v$ along with it. Note that $a_{i}$ always sees a right marking until $v_{b}$. Now when it moves to $v_{b}$, if it is not consumed it must see no such right marking as $a_{j}$ failed to mark it at round $t$. In this case $a_{i}$ leaves the node in clockwise direction to a new node. So here $a_{i}$ is able to detect the black hole. On the other hand if $a_{i}$ is consumed while it visits $v_{b}$, then it can not return back to $a_{k}$. When $a_{k}$ sees that $a_{i}$ has not returned it interprets that $a_{i}$ must have been consumed at the black hole which is the next node along clockwise direction. Thus in this case also at least one agent can detect the black hole.
@@
Case-II: $a_{j}$ was in state Backtrack at round $t$. In this situation, let without loss of generality, the node $v_{2}$ be the $home$ of $a_{i}$ and, it is also the cautious start node. Let $v_{1}$ be the farthest node along counter-clockwise direction in $Seg(a_{j})$ (i.e., the $home$ of $a_{j}$). Now, if there exists any round $t^{\prime}<t$ when $a_{j}$ was in state Forward, then it must have started this state from $v_{1}$ and moved in a clockwise direction until it reaches the node $v_{2}$ while erasing all the left markings from each node it traverses in $Seg(a_{j})$. This means that, when $a_{j}$ started the state Backtrack from $v_{2}$, there does not exist any node in $Seg(a_{j})$ with left markings. So before $a_{j}$ is consumed by the black hole node $v_{b}$ (say) at round $t$, it must have marked all the nodes with left, starting from the next counter-clockwise node of $v_{2}$ to the adjacent clockwise neighbor of $v_{b}$ (as before writing this message at $v_{b}$, the agent gets consumed by the black hole). Let without loss of generality, $a_{i}$ be the lowest ID agent among $a_{i}$ and $a_{k}$, hence it starts in state Cautious-Leader, whereas $a_{k}$ starts in state Cautious-Follower, at some round $t_{0}>t$. This means $a_{i}$ is the first agent to move alone in the next counter-clockwise neighbour say, $v$. After which, only if it sees a left message then only it moves back in clockwise direction at the node of $a_{k}$, and in the next round both these agents reach the node $v$. Observe, $a_{i}$ always finds a left message until the node $v_{b}$. Whenever it reaches $v_{b}$, either it gets consumed by the black hole, or it finds that no left marking is present at the current node. This triggers $a_{i}$ to detect the current node to be the black hole and moves in counter-clockwise direction to a new node. Otherwise, if it also gets consumed by the black hole, then in the next round it is unable to return to $a_{k}$, which triggers $a_{k}$ to conclude that $a_{i}$ must have been consumed by the black hole as well, and it correctly detects the black hole to be the next node in the counter-clockwise direction. Thus for each scenario there exists at least one agent which is able to correctly detect the black hole location.
@@
Note that within at most $3n$ number of rounds after both alive agent starts cautious walk from cautious start node, the black hole will be detected by at least one agent. So from Lemma 5.19 and Lemma 5.21 within atmost $10n$ rounds after the first agent is consumed by the black hole there exists atleast one agent that knows the exact location of the black hole which can now explore the ring $R$ perpetually avoiding the black hole. So if all three agents strats from three different nodes the PerpExploration-BBH will be solved if each node has a whiteborad of memory $O(\log n)$. Now for the case where three agents starts from two different nodes we have the following remark.

### v2

@@ 11.2 Correctness and Complexity
Case-II: Let $a_{j}$ gets destroyed by the Byzantine black hole at round $t$ while in state Backtrack. We have to show that the agent that changes the state to Gather must store the dir type message with direction component counter-clockwise (as any agent can only move in the counter-clockwise direction, in state Backtrack). It will be enough to show that the agent that changes its state to Gather must have changed it from state Forward (as only in this case the state changing agent stores the dir type message having a counter-clockwise direction component in its local memory). Let $t^{\prime}$ be the round when $a_{j}$ was in state Initial for the last time. We claim that $a_{i}$ and $a_{k}$ can not change to state Gather before round $t^{\prime}+3n+1$. Note that at round $t^{\prime}$ both $a_{i}$ and $a_{k}$ are at their corresponding $home$ in the state Initial. Now if any one of them changes to state Gather, the earliest it can happen is at round $t^{\prime}+n+1$ when both of them are in state Forward. In this case, the agent that changes to state Gather must have found a visited type message at the current node (i.e., for $a_{i}$ it is $home$ of $a_{k}$ and for $a_{k}$ it is $home$ of $a_{j}$) at round $t^{\prime}+n+1$. Note that during round $t^{\prime}$ any previous messages are erased from both $home$ of $a_{k}$ and $home$ of $a_{j}$ by $a_{k}$ and $a_{j}$, respectively, and no other agent can visit and alter data at these nodes before $t^{\prime}+n+1$ round. Hence none of $a_{i}$ and $a_{k}$ finds any visited type message at their current nodes at round $t^{\prime}+n+1$. Hence, none of these agents changes their state to Gather at round $t^{\prime}+n+1$. Next, they can only change their state to Gather at the round $t^{\prime}+3n+1$ when both of them (i.e., $a_{i}$ and $a_{k}$) are at their respective $home$. An agent among them changes its state to Gather at round $t^{\prime}+3n+1$ if it finds no visited type message at their current node (i.e., corresponding $home$). Note that $t^{\prime}+3n+1>t>t^{\prime}+n+1$ (as $a_{j}$ was in state Backtrack at round $t$). So, at the round $t^{\prime}+n+1$ all of $a_{i},a_{j}$ and $a_{k}$ are at nodes $home$ of $a_{k},home$ of $a_{i}$ and $home$ of $a_{j}$ in state Forward and writes a visited type message in the nodes, respectively. These messages can not be altered by any agent until round $t^{\prime}+3n+1$ (by a similar argument as in Case-I). Note that at round $t^{\prime}+3n+1$ both $a_{i}$ and $a_{k}$ are at their corresponding $home$ and both of them find a visited type message at these nodes left by $a_{j}$ and $a_{i}$ respectively (at round $t^{\prime}+n+1$). So, none of them changes to state Gather even at round $t^{\prime}+3n+1$. Next, they move into state Initial-Wait and wait at their $home$ until $t^{\prime}+4n$ round. Now, since no agent in state Initial-Wait can change its state to Gather2 until it meets with an agent in state Gather and all alive agents at round $t^{\prime}+3n+2$ are at state Initial-Wait all of them (i.e., $a_{i}$ and $a_{k}$) waits and changes state to Initial again at round $t^{\prime}+4n+1$. Next at round $t^{\prime}+4n+2$ both $a_{i}$ and $a_{k}$ erase all previous data at their corresponding $home$. and changes status to Forward again. Note that the next time $a_{i}$ and $a_{k}$ can change the state to gather must be at the round $t^{\prime}+5n+2$ when both of them are in state Forward, and are currently at the $home$ of $a_{k}$ and $home$ of $a_{j}$, respectively. Further, note that since $home$ of $a_{k}$ does not have any visited type message at round $t^{\prime}+5n+2$ (as $a_{k}$ erased any data at round $t^{\prime}+4n+2$ and no other agent can alter data there, after $t^{\prime}+4n+2$ and before $t^{\prime}+5n+2$), so $a_{i}$ can not change to state Gather at round $t^{\prime}+5n+2$. Also observe that the visited type message at the $home$ of $a_{j}$ written by $a_{k}$ during round $t^{\prime}+3n+1$ is still there at round $t^{\prime}+5n+2$ (after $t^{\prime}+3n+1$ before $t^{\prime}+4n+2$ since $a_{j}$ is already destroyed before $t^{\prime}+3n+1$, no agent is on $home$ of $a_{j}$ to erase the data, also after $t^{\prime}+4n+1$ till $t^{\prime}+5n+1$ only $a_{k}$ can visit $home$ of $a_{j}$ in state Forward and Back-Wait, but in these states it does not alter any data at the $home$ of $a_{j}$) $a_{k}$ finds it during the round $t^{\prime}+5n+2$ in state Forward and changes its state to Gather.
@@
Proof 11.8.
Suppose an agent $a_{i}$ in state Back-Wait gets a dir type message with clockwise direction component at some round $t$. This implies there exists another agent $a_{k}$ that has changed its state to Gather after storing a dir type message having a direction component clockwise at some round $t^{\prime}<t$. This can only happen if $a_{k}$ was in state Backtrack at the beginning of round $t^{\prime}$. So at the beginning of round $t^{\prime}$, $a_{i}$ was also in state Backtrack, and thus at round $t^{\prime}$, $a_{i}$ changes its state to Initial-Wait. Note that $a_{k}$ meets and shares dir type message with $a_{i}$ while $a_{i}$ is still at state Initial-Wait. This contradicts our assumption that $a_{i}$ gets dir message at state Back-Wait. Thus, if $a_{i}$ finds a dir type message in the state Back-Wait then it must have the direction component counter-clockwise.
With a similar argument, we can also prove the following lemma.
@@
Let $a_{i}$, $a_{j}$, and $a_{k}$ be three agents executing the algorithm PerpExplore-Scat-Whitbrd, and suppose $a_{j}$ be the first agent to enter the Byzantine black hole, while exploring the segment $Seg(a_{j})$. Let $v_{1}$ be the $home$ of $a_{j}$ and $v_{2}$ be the furthest node from $v_{1}$ along the clockwise direction which is inside $Seg(a_{j})$. We define the Cautious start node to be $v_{1}$ if $a_{j}$ is destroyed by the Byzantine black hole during state Forward. Otherwise, if $a_{j}$ is destroyed by the Byzantine black hole in state Backtrack then, $v_{2}$ is defined to be the Cautious start node.
@@
Let the first agent be destroyed by the Byzantine black hole at some round $t>0$, then there exists a round $t^{\prime}>t$, at which the remaining alive agents reach the cautious start node and change their state to Cautious-Leader and Cautious-Follower.
Proof 11.12.
@@
Let $a_{i}$, $a_{j}$ and $a_{k}$ be three agents, exploring $Seg(a_{i})$, $Seg(a_{j})$ and $Seg(a_{k})$, respectively, where $Seg(a_{i})\cap Seg(a_{j})=home$ of $a_{i}$, $Seg(a_{i})\cap Seg(a_{k})=home$ of $a_{k}$ and $Seg(a_{k})\cap Seg(a_{j})=home$ of $a_{j}$. Let $a_{j}$ be the agent that gets destroyed by the Byzantine black hole while exploring $Seg(a_{j})$, and then we have the following cases:
@@
Case-I: $a_{j}$ is destroyed by the Byzantine black hole while it is moving in a clockwise direction, i.e., in state Forward at round $t>0$. Note that in this case, the cautious start node is the $home$ of $a_{j}$. This implies there exists a round $0<t_{0}<t<t_{0}+n$, where $a_{j}$ was in state Initial. Note that, $a_{j}$ can not write any visited type message at $home$ of $a_{i}$, as it gets destroyed before reaching that node. So, at round, $t_{0}+3n+1$, when $a_{i}$ is at its own $home$ in state Backtrack and checks any visited type message, it finds none exists. This triggers $a_{i}$ to change its state to Gather from Backtrack with the corresponding dir = (clockwise, NULL) (which is indeed the correct direction, refer Lemma 11.5). Further, the agent finds $a_{k}$, which is currently waiting at its own $home$ in state Initial-Wait (as it does not find any anomaly, so from Backtrack it changed to state Initial-Wait at round $t_{0}+3n+1$). So, the moment $a_{i}$ reaches the $home$ of $a_{k}$, it takes one additional round to store the dir type message at the current node and then changes its state to Gather1. On the other hand, whenever $a_{k}$ finds a dir type message is written at its current node, it also changes its state to Gather1, i.e., at round $t_{0}+3n+1<t^{\prime\prime}<t_{0}+4n$ both agents change their state to Gather1. After which, they together start moving in a clockwise direction (11.1), until they reach a node which has a home type message with the ID component, different from the IDs of $a_{i}$ and $a_{k}$. Note that this must be the $home$ of $a_{j}$, as the home type message $a_{j}$ has written at round $t_{0}$, cannot be erased by any other agent except $a_{j}$. Also $a_{j}$ can only erase this in state Initial at round $t_{0}+4n+2$ if it would have returned back, but since it is already destroyed between round $t_{0}+1$ and $t_{0}+n-1$, hence this possibility never arises, so the home type message remains, when $a_{i}$ and $a_{k}$ together reaches this node. After which they change their state to Cautious-Leader and Cautious-Follower depending on their IDs.
@@
Case-II: $a_{j}$ is destroyed by the Byzantine black hole while it is moving in a counter-clockwise direction, i.e., in state Backtrack at round $t>0$. Note that in this case, the cautious start node is the $home$ of $a_{i}$. Let $t_{0}$ be the round when $a_{j}$ was in state Initial the last time. This implies $t_{0}+2n<t<t_{0}+3n$, now by similar argument as explained in Case-II of Lemma 11.5, $a_{k}$ changes its state to Gather while storing the dir = (counter-clockwise, NULL) message, at the $home$ of $a_{j}$ from state Forward, and starts moving in a counter-clockwise direction. Note that at round $t_{0}+5n+2$ as $a_{i}$ did not find any anomaly, so it changes its state to Back-Wait from Forward. Hence, at round $t_{1}$, where $t_{1}<t_{0}+6n$, $a_{k}$ finds $a_{i}$, while $a_{i}$ is still in state Back-Wait. This triggers $a_{k}$ to change its state to Gather2 at round $t_{1}$ while updating the dir type message to (Counter-clockwise, ID’), where ID’ is the ID of $a_{i}$. Then at the same round, $a_{k}$ writes the updated message at the current node (i.e., $home$ of $a_{k}$). Whenever $a_{i}$ sees this dir type message (at round $t_{1}+1$) it also changes its state to Gather2. Next in state Gather2, both start to move counter-clockwise (from round $t_{1}+2$) and continue to move until they find a home type message with ID matching the ID of the dir type message. Note that, this node is nothing but the $home$ of $a_{i}$ (as the ID component of dir type message stores the ID of $a_{i}$). Note that at round $t_{0}+4n+2$, $a_{i}$ written a home type message at its own $home$. This message can be erased only again at round $t_{0}+8n+2$ (i.e., when $a_{i}$ reaches its $home$ again in state Initial). But since $a_{i}$ changed its state to Gather2 before $t_{0}+6n+1$ it cannot move back to state Initial again. So,when $a_{k}$ and $a_{i}$ reaches $home$ of $a_{i}$, they finds the home type message. Hence, both $a_{k}$ and $a_{i}$ reach the cautious start node within $t_{0}+7n$ and further change their states to Cautious-Leader and Cautious-Follower, depending on their IDs.
@@
Let $a_{i}$ and $a_{k}$ be the two agents that start the states Cautious-Leader and Cautious-Follower from the cautious start node, then within finite rounds of executing algorithm PerpExplore-Scat-WhitBrd, at least one agent detects the location of the Byzantine black hole.
Proof 11.14.
@@
Let $a_{j}$ be the agent that has been destroyed by the Byzantine black hole at round $t>0$. Now there are two cases based on the state of $a_{j}$ at round $t$.
@@
Case-I: $a_{j}$ was in state Forward at round $t$. In that case, the node $v_{1}=home$ of $a_{j}$ is the cautious start node. Let $v_{2}$ be the farthest node along the clockwise direction in $Seg(a_{j})$ (i.e., the clockwise nearest $home$ of $a_{j}$). Now if there exists any round $t^{\prime}<t$ when $a_{j}$ was in the state Backtrack then it must have started its state Backtrack from $v_{2}$ and moved counter-clockwise until $v_{1}$ while erasing all right markings from each node in $Seg(a_{j})$. So when $a_{j}$ started the state Forward at its $home$ before being destroyed, all nodes in $Seg(a_{j})$ are without any right marking. This case will happen also when there is no round $t^{\prime}<t$ where $a_{j}$ was in state Backtrack, i.e., there does not exist any round before $t^{\prime}$, when $a_{j}$ was in state Forward. So before $a_{j}$ is destroyed at the Byzantine black hole, say $v_{b}$, at round $t$, it has marked right at all nodes starting from the next node of $v_{1}$ in the clockwise direction, up to the node just before $v_{b}$ ($v_{b}$ can not be marked as before marking it the agent $a_{j}$ is destroyed there). Let without loss of generality, $a_{i}$ is in state Cautious-Leader and $a_{j}$ is in state Cautious-Follower at $v_{1}$ at some round $t_{0}>t$. Then $a_{i}$ always moves ahead alone in the clockwise direction to a new node $v$.Then it moves back only if sees the right marking and brings $a_{k}$ to $v$ along with it. Note that $a_{i}$ always sees a right marking until $v_{b}$. Now when it moves to $v_{b}$, if it is not destroyed it must see no such right marking as $a_{j}$ failed to mark it at round $t$. In this case, $a_{i}$ leaves the node in the clockwise direction to a new node. So here $a_{i}$ is able to detect the Byzantine black hole. On the other hand, if $a_{i}$ is destroyed while it visits $v_{b}$, then it can not return back to $a_{k}$. When $a_{k}$ sees that $a_{i}$ has not returned it interprets that $a_{i}$ must have been destroyed at the Byzantine black hole which is the next node along the clockwise direction. Thus in this case also at least one agent can detect the Byzantine black hole.
@@
Case-II: $a_{j}$ was in state Backtrack at round $t$. In this situation, let without loss of generality, the node $v_{2}$ be the $home$ of $a_{i}$ and, it is also the cautious start node. Let $v_{1}$ be the farthest node along counter-clockwise direction in $Seg(a_{j})$ (i.e., the $home$ of $a_{j}$). Now, if there exists any round $t^{\prime}<t$ when $a_{j}$ was in state Forward, then it must have started this state from $v_{1}$ and moved in a clockwise direction until it reaches the node $v_{2}$ while erasing all the left markings from each node it traverses in $Seg(a_{j})$. This means that, when $a_{j}$ started the state Backtrack from $v_{2}$, there does not exist any node in $Seg(a_{j})$ with left markings. So before $a_{j}$ is destroyed by the Byzantine black hole node $v_{b}$ (say) at round $t$, it must have marked all the nodes with left, starting from the next counter-clockwise node of $v_{2}$ to the adjacent clockwise neighbor of $v_{b}$ (as before writing this message at $v_{b}$, the agent gets destroyed by the Byzantine black hole). Let without loss of generality, $a_{i}$ be the lowest ID agent among $a_{i}$ and $a_{k}$, hence it starts in state Cautious-Leader, whereas $a_{k}$ starts in state Cautious-Follower, at some round $t_{0}>t$. This means $a_{i}$ is the first agent to move alone in the next counter-clockwise neighbor say, $v$. After which, only if it sees a left message then only it moves back in the clockwise direction at the node of $a_{k}$, and in the next round both these agents reach the node $v$. Observe that, $a_{i}$ always finds a left message until the node $v_{b}$. Whenever it reaches $v_{b}$, either it gets destroyed by the Byzantine black hole, or it finds that no left marking is present at the current node. This triggers $a_{i}$ to detect the current node to be the Byzantine black hole and moves in the counter-clockwise direction to a new node. Otherwise, if it also gets destroyed by the Byzantine black hole, then in the next round it is unable to return to $a_{k}$, which triggers $a_{k}$ to conclude that $a_{i}$ must have been destroyed by the Byzantine black hole as well, and it correctly detects the Byzantine black hole to be the next node in the counter-clockwise direction. Thus for each scenario, there exists at least one agent that is able to correctly detect the Byzantine black hole location.
@@
Note that within at most $3n$ number of rounds after both alive agent starts cautious walk from the cautious start node, the Byzantine black hole will be detected by at least one agent. So from Lemma 11.11 and Lemma 11.13 within at most $10n$ rounds after the first agent is destroyed by the Byzantine black hole there exists at least one agent that knows the exact location of the Byzantine black hole which can now explore the ring $R$ perpetually avoiding the Byzantine black hole. So if all three agents start from three different nodes the PerpExploration-BBH will be solved if each node has a whiteboard of memory $O(\log n)$. This proves the Theorem 5.4. We state the theorem again here for convenience.
Statement of theorem 5.4: Algorithm PerpExplore-Scat-Whitbrd solves
PerpExploration-BBH problem of a ring $R$ with $n$ nodes and with 3 synchronous agents initially scattered under the whiteboard model of communication
11.3 Modification to the algorithm to include the cases where the initial starting nodes can have multiplicity